OpenStudy (anonymous):

all things being equal, 10 factory workers laboring 10 hours per day produce 10 units of work in 10 days. how many of these units can be produced by 5 factory workers laboring 5 hours a day for 5 days

OpenStudy (anonymous):

@Hero here it is

OpenStudy (anonymous):

my teach told me the answer is 5/4 but i dont know how she arrived to it.

OpenStudy (anonymous):

*teacher

hero (hero):

Hold on a minute

OpenStudy (anonymous):

okay

hero (hero):

Okay, I have it.

hero (hero):

Basically, we can use the following proportion to solve this: \[\frac{m_1d_1h_1}{w_1} = \frac{m_2d_2h_2}{w_2}\] Where m = number of factory workers d = number of days h = number of hours w = units of work

OpenStudy (anonymous):

okay, so now we can plug in #s

hero (hero):

On the left side of the equation, we put the corresponding values for the first case. On the right side of the equation, we put the corresponding values for the second case. Do you see how we might set up the proportion at this point?

OpenStudy (anonymous):

yes, 10^3/10 = 5^3/w

hero (hero):

Nice job

OpenStudy (anonymous):

and.... w = 5/4

OpenStudy (anonymous):

Awesome! Thanks for all of your help, I really appreciate it. You really are a "Hero" !

hero (hero):

You're welcome